Assume the Black-Scholes framework. The 1-year futures price for stock LMN is $270. The volatility is...

  1. Assume the Black-Scholes framework. The 1-year futures price for stock LMN is $270. The volatility is 30%, and the interest rate is 4%. What is the price of a 280-strike call option price on the LMN futures contract, expiring 9 months from today?
Answer #1

We use Black-Scholes Model to calculate the price of the call option.

The price of a call option is:

C = (S0 * N(d1)) - (Ke-rt * N(d2))

where :

S0 = current spot price

K = strike price

N(x) is the cumulative normal distribution function

r = risk-free interest rate. This is 4%, or 0.04.

t is the time to expiry in years. This is (9/12), or 0.75.

d1 = (ln(S0 / K) + (r + σ2/2)*T) / σ√T

d2 = d1 - σ√T

σ = standard deviation of underlying stock returns

First, we calculate d1 and d2 as below :

  • ln(S0 / K) = ln(270 / 280). We input the same formula into Excel, i.e. =LN(270 / 280)
  • (r + σ2/2)*T = (0.04 + (0.302/2)*0.75
  • σ√T = 0.30 * √0.75

d1 = 0.1054

d2 = -0.1544

N(d1) and N(d2) are calculated in Excel using the NORMSDIST function and inputting the value of d1 and d2 into the function.

N(d1) = 0.5420

N(d2) = 0.4386

Now, we calculate the price of the call option as below:

C = (S0 * N(d1))   - (Ke-rt * N(d2)), which is (270 * 0.5420) - (280 * e(-0.04 * 0.75))*(0.4386)    ==> $27.1416

Price of call option is $27.1416